# Relativistic unitary coupled cluster method for ground-state molecular properties

### v1

@@ Abstract
We propose a relativistic unitary coupled cluster (UCC) expectation value approach for computing first-order properties of heavy-element systems. Both perturbative (UCC3) and non-perturbative (qUCC) commutator-based formulations are applied to evaluate ground-state properties, including the permanent dipole moment (PDM), magnetic hyperfine structure (HFS) constant, and electric field gradient (EFG). The results are compared with available experimental data and those from conventional coupled cluster (CC) calculations. The non-perturbative commutator-based approach truncated at the singles and doubles level (qUCCSD) exhibits markedly better agreement with both CCSD and experiment than the perturbative UCC3 method, likely due to its improved treatment of relaxation effects.
@@ I Introduction
A comprehensive understanding of chemical processes requires consideration of not only molecular energetics but also intrinsic properties such as dipole moments, hyperfine coupling constants, and other measurable parameters. Therefore, quantum chemical calculations of atomic and molecular properties have become a fundamental aspect of modern-day computational chemistry.[1, 2, 3, 4] However, accurate calculation of properties of atoms and molecules containing heavy elements is highly challenging due to the necessity of incorporating both relativistic and electron correlation effects in a balanced manner, especially for systems where relativistic effects are significant.[5] One of the most effective ways to account for relativistic effects in a quantum mechanical calculation for many-electronic systems is through the Dirac–Hartree–Fock (DHF) method.[6, 7] The DHF method cannot take care of the dynamic interaction of opposite-spin electrons, which is known as the electron correlation effect. Among the various electron correlation methods available, the single-reference coupled cluster (CC) method [8, 9, 10, 11, 12, 13] has emerged as one of the most accurate and systematically improvable ones. The property calculation within the CC method can be performed using two alternative approaches: one is using an expectation value approach,[14, 15] and the second is using an analytic derivative technique.[16, 17, 18, 19] It should be noted that the two approaches yield different results even for first-order properties, except in the full CC limit.[14] In the domain of relativistic coupled cluster, the calculation of first-order property has been reported[20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] using both analytic derivative and expectation value approach. The analytic derivative approach is generally more advantageous for geometrical derivatives and higher-order properties.[19] The calculation of second and higher-order properties with the framework of the relativistic CC method has been recently reported.[31, 32, 33, 34] In addition to the standard formulation of the coupled cluster method, there exist alternate ansatzes within the CC framework.[35] One can use a unitary ansatz[36] to arrive at a hermitian formulation of coupled cluster.[37, 38, 39] In recent times, the unitary coupled cluster(UCC) method has gained significant attention, not only for its computational advantages but also because of its potential applications in quantum computing.[40, 41, 42, 42] The UCC method has a particular advantage: the first-order properties calculated using both expectation value and analytic derivative will yield the same answer for a particular truncation of the cluster operator.[38] Moreover, because the standard CC energy functional is non-Hermitian, calculating first-order ground-state properties with the analytic derivative technique requires not only the CC amplitudes but also an additional set of left-vector amplitudes to ensure that the energy functional is stationary. In the UCC expectation value formalism, one can obtain the first-order properties with a single set of amplitudes. However, unlike the conventional CC theory, the UCC framework does not allow for a natural truncation of the similarity-transformed Hamiltonian. As a result, an imposed truncation is necessary. Any arbitrary truncation of the UCC functional may compromise the size-extensivity of the energy.[38] To address this challenge, Bartlett and co-workers[43] proposed a truncation scheme based on perturbative analysis of the UCC energy functional, preserving the size-extensivity of the energy. Building on this, Taube and Bartlett[44] introduced an improved truncation scheme that yields exact results for two-electron systems. However, for larger and more complex molecules, the perturbative approximation to the UCC theory(UCC(n))often fails due to the poor convergence behavior of the underlying low-order Møller-Plesset(MP) series.[45] To overcome this limitation, recent efforts have focused on truncating the UCC expansion based on the rank of the nested commutators[46, 47, 48] using Bernoulli expansion, which has shown improved performance in systems with irregular MP convergence. Furthermore, Liu et al.[48] proposed the UCCn method, which applies a perturbative truncation of the Bernoulli expansion and offers enhanced accuracy in such challenging cases. The UCC3 method has been particularly well studied[48, 49, 50, 51] in recent times for ground and excited states of small molecules. Recent work by DePrince and co-workers[45] demonstrated that the commutator-based truncation scheme achieves faster convergence toward the Full CI limit and provides more reliable results than the perturbative truncation-based approximations, especially for molecular systems away from their equilibrium geometries. In spite of many attractive features, the report of the property calculation using UCC is rather limited. Bartlett and co-workers have reported an analytic gradient technique for the UCC(4) method.[52] Sur et al. have used the relativistic UCC3 method to calculate nuclear quadrupole moment, hyperfine constants, and transition properties of atoms.[46] Dreuw and co-workers[50] have recently reported ground and excited state properties using the UCC3 method. One of the main issues with property calculations using low-order Møller–Plesset (MP) theory UCC methods is that they yield less accurate results than the standard CC approach.[52] The non-perturbative quadratic unitary coupled cluster method (qUCC) of Cheng and co-workers has been shown to give comparable performance to that of the standard CC method for energy calculations.[47] We have recently reported a relativistic variant of the qUCC method.[53] The aim of this paper is to extend the relativistic UCC method for the first-order property calculation using the expectation value approach. The structure of this paper is as follows: Section II presents the theoretical framework of the method, Section III provides the computational details, and Section IV presents the results of molecular property calculations for selected systems. Finally, conclusions are summarized in Section V.
@@ II.1 Relativistic Unitary Coupled Cluster Theory
The DHF method is generally the starting point for all relativistic single-reference electron correlation calculations. It extends the Hartree-Fock method by incorporating the principle of special relativity via the Dirac-Coulomb (DC) Hamiltonian,[7] , and for a molecular system under the Born-Oppenheimer approximation, can be defined as,
where denotes the momentum operator, is the rest mass, and is the speed of light. The operator represents the potential energy of the -th electron in the field of the nucleus A with a Gaussian charge distribution. The symbols and are the Dirac matrices, and denotes the identity matrix. The DHF method can be represented in matrix form
In this representation, and correspond to the large and small components of the four-component wavefunction, each of the components taking the form of a two-spinor. The operator represents the nuclear-electron interaction. Whereas and represent the Coulomb and exchange operators, respectively. The Pauli spin matrices are denoted by .
The Eq. (1) can be rewritten in the occupation number representation as
Where is the Fock matrix, defined as
@@
Here, the indices p, q, r, s represent positive-energy four-component spinors within the framework of the no-pair approximation [54]. To account for the effects of electron correlation, we employ the UCC ansatz on top of the DHF reference . In the UCC method, the ground-state wave function is expressed as
The is the excitation operator, while its Hermitian conjugate accounts for the corresponding de-excitation cluster operator. The difference forms an anti-Hermitian operator, i.e, , ensuring that the exponential operator is unitary and thus preserves the norm of the wave function. The cluster operator in the singles and doubles truncation is given by
where , , , , and , , , symbols are used to denote the occupied and virtual spinors, respectively. The ground state energy is obtained as
The cluster amplitudes in the projection based unitary coupled cluster are determined by projecting the electronic Schrödinger equation on to the excited state determinants as,
is the similarity-transformed Hamiltonian defined as,
@@
One major challenge in UCC theory is that, unlike traditional CC theory, there is no natural finite truncation of the Baker–Campbell–Hausdorff (BCH) expansion for . This arises due to the presence of both excitation and de-excitation operator cluster operators in . Therefore, the expansion requires an imposed truncation. One needs to take special precautions to ensure the size extensivity of the truncated unitary coupled cluster method. This can be done either by applying a perturbative truncation[43] based on Møller-Plesset (MP) perturbation theory or by using a commutator-based approach[48]. The commutator-based truncation schemes using the Bernoulli numbers [55, 56] have been found to be particularly effective. It is generally derived using the superoperator (denoted by a double hat) approach[57]. A superoperator (), when applied on any arbitrary operator (), gives rise to a commutator
Using the above definition, the Eq. (12) can be written as
The above expression is equivalent to the BCH expansion Eq. (13).
Now, one can separate the using the superoperator
Where denotes the exponential Taylor series expressed as
The inverse of the above function can be expressed as
The denotes the Bernouli numbers
By left multiplying Eq. (17) with , one can arrive at
One can obtain expressions for the iterative generation of as
@@
As it can be seen that the iterative generation of only includes a singly nested commutator in , which makes it advantageous over other truncation schemes to UCC. To use the iterative rule to generate the , the can be split into two parts
Where “N” denotes the non-diagonal part of the operator, which involves all pure excitation and de-excitation operators, and “R” represents the rest of the part obtained after excluding the non-diagonal part. With this definition, the UCC amplitude equations can be rewritten as
Assuming the Fock operator to be block diagonal. The iterative formula for the similarity-transformed fluctuation potential can now be written as
The above expression allows the construction of the only from the knowledge of the rest of the part of the similarity-transformed fluctuation potential. One can start with the guess and put into Eq. (26) to obtain
Which can then be plugged into Eq. (26) to obtain and so on. Therefore, the total can be expressed as
@@
and so on. One can derive non-perturbative approximations to the UCC method by including different orders of in the energy and amplitude equations. Among the various approximations that the above approach offers, the qUCC method offers the best compromise between computational cost and accuracy[45]. The qUCC energy and amplitudes are as follows
The qUCC method is generally used in singles and doubles approximation (qUCCSD)[47]. Extension of the qUCCSD method to non-iterative triples correction has also been achieved[58]. Alternatively, one can derive a perturbative approximation to UCC by truncating the amplitudes in Eq. (25) in perturbation order, and it is generally denoted UCCn, where n denotes the order of perturbation. It is important to note that the UCCn approach differs from Bartlett’s UCC(n) framework[43], where the ground state energy is accurate through the order n in MP perturbation theory. In the present formulation, the amplitude equations are consistent through the n-th order.
II.2 First order Property Calculation using UCC Expectation Value Formalism
We have formulated the property calculation within the UCC method using an ADC-like intermediate state representation[59, 50]. The wave function for the Ith excited state can be defined as
Where is a CI-like excitation operator for the ground state, is a unity operator, and the ground state expectation value corresponding to an arbitrary operator () can be written as
Consequently, one can define the ground-state first-order property at the qUCCSD level
Subsequently, the excited state first-order property can be represented as
Where will result in the first-order property corresponding to excited state, and corresponds to the transition property due to excitation from to excited state. The programmable expressions for the qUCCSD first-order ground state property calculation are provided in the supporting information. One can derive the corresponding UCC3 expression by neglecting terms that contain amplitude products beyond fourth order in perturbation from Eq. (38). To assess the performance of the new UCC expectation value approach, we have calculated the permanent dipole moment (PDM), magnetic hyperfine structure (HFS) constant, and electric field gradient (EFG), and compared them with the CCSD Z-vector method and experimental results wherever available.
@@ II.2.1 Permanent Dipole Moment (PDM)
The PDM operator () of a molecular system can be expressed as the sum of the electronic contribution (EC) and the nuclear contribution (NC),
In the above equation, e is the charge of the electron, is the position vector of the electron from the origin, is the atomic number and is the position vector of nucleous. Now eq (37) can be re-written as
The electronic contribution to in Eq. (41) depends on the choice of correlation method and basis set, whereas the nuclear part is independent of both the correlation method and the basis set; it remains constant for a particular molecule with a certain charge and at a specific bond length.
@@ II.2.2 Hyperfine Structure (HFS) Constant
The magnetic hyperfine splitting in atoms arises from the interaction between the electromagnetic field generated by the electrons and the magnetic dipole moment of the nucleus. Within the magnetic dipole approximation, the vector potential at the position of electron due to the nuclear magnetic moment is given by
where is the position vector of the electron relative to the nucleus.In a relativistic framework, the corresponding magnetic hyperfine interaction for all electron can be written as
where denotes the Dirac matrices of electron . For a diatomic molecule, the magnetic hyperfine interaction can be separated into parallel and perpendicular components of the magnetic HFS constant. The z projection of the expectation value of the corresponding HFS operator gives the while the x/y projections give the constant, expressed as
Where denotes the nuclear spin quantum number and is the projection of total electronic angular momentum along the z axes of the diatomic molecule.
@@ II.2.3 Electric Field Gradient (EFG)
The EFG is a traceless symmetric second-rank tensor that describes the spatial variation of the electric field around a nucleus due to the distribution of surrounding charges. The EFG provides valuable information about the local electronic structure and symmetry of atoms or molecules, particularly in cases where the environment of the nucleus is asymmetric. Within the principal axes coordinate system, the expectation value of the component of EFG at the position () of nucleus K is defined as
The expectation value of the above operator will lead to two terms
The first term in the above equation is the electronic contribution, named , which depends on the ground state wave-function. The second term , which arises from the other nuclei in the molecule, depends only on the position of that remaining nucleus and is independent of the electronic wave function.

### v2

@@ Abstract
We propose a relativistic unitary coupled cluster (UCC) expectation-value approach for computing first-order properties of heavy-element systems. Both perturbative and non-perturbative commutator-based formulations are applied to evaluate ground-state properties, including the permanent dipole moment (PDM), magnetic hyperfine structure (HFS) constant, and electric field gradient (EFG). The results are compared with available experimental data and with those obtained using conventional coupled cluster (CC) calculations. The non-perturbative commutator-based approach truncated at the singles-and-doubles level exhibits markedly better agreement with both the coupled cluster Z-vector method and experimental values than the perturbative variant, likely due to its improved treatment of relaxation effects.
@@ I Introduction
A comprehensive understanding of chemical processes requires consideration of not only molecular energetics but also intrinsic properties such as dipole moments, hyperfine coupling constants, and other measurable parameters. Therefore, quantum chemical calculations of atomic and molecular properties have become a fundamental aspect of modern computational chemistry.[1, 2, 3, 4] However, accurate calculation of properties of atoms and molecules containing heavy elements is highly challenging due to the necessity of incorporating both relativistic and electron correlation effects in a balanced manner, especially for systems in which relativistic effects are significant.[5] One of the most effective ways to account for relativistic effects in a quantum mechanical calculation for many-electron systems is through the Dirac–Hartree–Fock (DHF) method.[6, 7] The DHF method does not account for the dynamic interaction of opposite-spin electrons, which is known as the electron correlation. Among the various electron correlation methods available, the single-reference coupled cluster (CC) method [8, 9, 10, 11, 12, 13] has emerged as one of the most accurate and systematically improvable approaches. The property calculation within the CC method can be performed using two alternative approaches: an expectation value approach,[14, 15] and an analytic derivative technique.[16, 17, 18, 19] It should be noted that the two approaches yield different results even for first-order properties, except in the full CC limit.[14] In the domain of relativistic CC method, the calculation of first-order properties has been reported[20, 5, 21, 22, 23, 24, 25, 26, 27, 28, 29] using both analytic derivative and expectation value approaches. The analytic derivative approach is generally more advantageous for geometric derivatives and higher-order properties.[19] The calculation of second and higher-order properties within the framework of the relativistic CC method has been reported recently.[30, 31, 32, 33] In addition to the standard formulation of the CC method, alternate ansatz exist within the CC framework.[34] One can use a unitary ansatz[35] to arrive at a Hermitian formulation of coupled cluster theory.[36, 37, 38] In recent years, the unitary coupled cluster (UCC) method has gained significant attention, not only for its computational advantages but also because of its potential applications in quantum computing.[39, 40, 41, 41] The UCC method has a particular advantage: first-order properties calculated using both expectation-value and analytic-derivative approaches yield the same result for a given truncation of the cluster operator.[37] Moreover, because the standard CC energy functional is non-Hermitian, the calculation of first-order ground-state properties with the analytic derivative technique requires not only the CC amplitudes but also an additional set of left-vector amplitudes to ensure that the energy functional is stationary. In the UCC expectation value formalism, first-order properties can be obtained using a single set of amplitudes. However, unlike conventional CC theory, the UCC framework does not allow a natural truncation of the similarity-transformed Hamiltonian. As a result, an explicit truncation is required. Any arbitrary truncation of the UCC functional may compromise the size extensivity of the energy.[37] To address this challenge, Bartlett and co-workers[42] proposed a truncation scheme based on perturbative analysis of the UCC energy functional, thereby preserving size extensivity. Building on this work, Taube and Bartlett introduced an improved truncation scheme that yields exact results for two-electron systems.[43] However, for larger and more complex molecules, the perturbative approximation to the UCC theory [UCC(n)] often fails due to the poor convergence behavior of the underlying low-order Møller-Plesset(MP) series.[44] To overcome this limitation, recent efforts have focused on truncating the UCC expansion based on the rank of nested commutators[45, 46, 47] using the Bernoulli expansion, which has shown improved performance in systems with irregular MP convergence. Furthermore, Liu et al.[47] proposed the UCCn method, which applies a perturbative truncation of the Bernoulli expansion and offers an alternate hierarchy of perturbative approximations to UCC methods. The UCC3 method has been particularly well studied[47, 48, 49, 50] in recent years for ground and excited states of small molecules. Recent work by DePrince and co-workers[44] demonstrated that the commutator-based truncation scheme achieves faster convergence toward the full configuration interaction (FCI) limit and provides more reliable results than perturbative truncation-based approximations, especially for molecular systems away from their equilibrium geometries. Despite its many attractive features, reports of property calculations using UCC remain limited. Bartlett and co-workers have reported an analytic gradient technique for the UCC(4) method.[51] Sur et al. have used the relativistic UCC3 method to calculate nuclear quadrupole moments, hyperfine constants, and transition properties of atoms.[45] Dreuw and co-workers[49] have recently reported ground- and excited-state properties using the UCC3 method. One of the main issues with property calculations using low-order Møller–Plesset (MP)-based UCC methods is that they yield less accurate results than the standard CC approach.[51] The non-perturbative quadratic unitary coupled cluster (qUCC) developed by Cheng and co-workers has been shown to provide performance comparable to that of the standard CC method for energy calculations.[46] We have recently reported a relativistic variant of the qUCC method.[52] This paper aims to perform first-order property calculations using the expectation-value approach. The structure of this paper is as follows: Section II presents the theoretical framework, Section III provides the computational details, and Section IV presents the results of molecular property calculations for selected systems. Finally, conclusions are summarized in Section V.
@@ II.1 Relativistic Unitary Coupled Cluster Theory
The DHF method is generally the starting point for all relativistic single-reference electron correlation calculations. It extends the Hartree-Fock method by incorporating the special theory of relativity via the Dirac-Coulomb (DC) Hamiltonian [7] and for a molecular system, under the Born-Oppenheimer approximation, can be defined as
where denotes the momentum operator, is the rest mass, and is the speed of light. The operator represents the potential energy of the -th electron in the field of the nucleus A. The symbols and are the Dirac matrices, and denotes the identity matrix. The DHF equations can be represented in matrix form as
In this representation, and correspond to the large and small components of the four-component (4c) wavefunction, each of the components taking the form of a two-spinor. The operator represents the nuclear-electron interaction. The operators and represent the Coulomb and exchange operators, respectively. The Pauli spin matrices are denoted by .
Equation (1) can be rewritten in the occupation number representation as
Where is the Fock operator, defined as
@@
Here, the indices p, q, r, and s represent positive-energy four-component spinors within the framework of the no-pair approximation [53]. To account for electron-correlation effects, we employ the UCC ansatz on top of the DHF reference state . In the UCC method, the ground-state wave function is expressed as
The is the excitation operator, while its Hermitian conjugate , accounts for the corresponding de-excitation cluster operator. The difference forms an anti-Hermitian operator, i.e, , ensuring that the exponential operator is unitary and thus preserves the norm of the wave function. Within the singles and doubles truncation, the cluster operator is given by
where , , , , and , , , denote the occupied and virtual spinors, respectively. The ground-state energy is obtained as
The cluster amplitudes in the projection based UCC are determined by projecting the electronic Schrödinger equation onto the excited state determinants as,
Here, is the similarity-transformed Hamiltonian, defined as
@@
One major challenge in the UCC theory is that, unlike traditional CC theory, there is no natural finite truncation of the BCH expansion for . This arises due to the presence of both excitation and de-excitation operators in . Therefore, the expansion requires an explicit truncation. Special care must be taken to ensure the size extensivity of the truncated unitary coupled cluster method. This can be achieved either by applying a perturbative truncation[42] based on Møller-Plesset (MP) perturbation theory or by using a commutator-based approach[47]. The commutator-based truncation schemes employing Bernoulli numbers [54, 55] have been found to be particularly effective. These schemes are generally derived using the superoperator (denoted by a double hat) formalism.[56] A superoperator (), when applied to an arbitrary operator (), gives rise to a commutator
Using the above definition, Eq. (12) can be written as
The above expression is equivalent to the BCH expansion given in Eq. (13).
The operator can now be separated using the superoperator formalism as
Here, denotes the exponential Taylor series expressed as
The inverse of this function can be expressed as
Here, denotes the Bernoulli numbers
By left-multiplying Eq. (17) by , one obtains
Expressions for the iterative generation of is
@@
It can be seen that the iterative generation of includes only a singly nested commutator in , which makes it advantageous compared to other UCC truncation schemes. To apply the iterative procedure for generating , the operator is partitioned into two parts:
Here, “N” denotes the non-diagonal part of the operator, which contains all pure excitation and de-excitation operators, while “R” represents the remaining part obtained after excluding the non-diagonal part. With this definition, the UCC amplitude equations can be rewritten as
The Fock operator is assumed to be block diagonal. The iterative formula for the similarity-transformed fluctuation potential can now be written as
The above expression allows the construction of the only from knowledge of the rest of the part of the similarity-transformed fluctuation potential. One can start with the guess and put into Eq. (26) to obtain
Which can then be plugged into Eq. (26) to obtain and so on. Therefore, the total similarity-transformed Hamiltonian can be expressed as
@@
and so on. Non-perturbative approximations to the UCC method can be derived by including different orders of in the energy and amplitude equations. Among the various approximations offered by this approach, the qUCC method provides the best compromise between computational cost and accuracy.[44] The qUCC energy and amplitudes are given by
The qUCC method is generally employed within the singles-and-doubles approximation (qUCCSD).[46] Extensions of the qUCCSD method to include noniterative triples corrections have also been developed.[57] Alternatively, perturbative approximations to UCC can be derived by truncating the amplitudes in Eq. (25) according to perturbation order, and it is generally denoted UCCn, where n denotes the order of perturbation. It is important to note that the UCCn approach differs from Bartlett’s UCC(n) framework[42], where the ground state energy is accurate through the order n in MP perturbation theory. In the present formulation, the amplitude equations are consistent through the n perturbation order.
II.2 First order Property Calculation using the UCC Expectation-Value Formalism
The wave function for the Ith excited state in the UCC formalism is defined as
Here, is a CI-like excitation operator. For the ground state, is an identity operator. The corresponding energy is defined as
The first derivative of the energy with respect to an arbitrary perturbation () can be written as
All possible Slater determinants can be partitioned into a principal space () and a complementary space ().
such that all () determinants spans the basis of .
In the FCI limit, the above expression can be rewritten (see the Supporting Information for details) as
However, for the truncated space, this expression is only an approximation. Accordingly, the first-order property corresponding to an arbitrary operator () in the qUCCSD expectation-value formalism can be written as
When , the expression yields the first-order property corresponding to the excited state, whereas corresponds to a transition property between the and excited states. Consequently, the ground-state first-order property at the qUCCSD level can be defined as
The above described approach for the first-order property calculation is similar in spirit to the ADC intermediate state representation[58, 49]. The programmable expressions for the UCC3 and qUCCSD first-order ground state property calculations are provided in the Supporting Information. The corresponding UCC3 expression can be derived by neglecting terms containing amplitude products beyond fourth order in perturbation theory in Eq.(42). To assess the performance of the new relativistic UCC3 and qUCCSD expectation-value approaches, we calculated the permanent dipole moment (PDM), magnetic hyperfine structure (HFS) constant, and electric field gradient (EFG) of small molecules containing heavy elements and compared the results with those obtained using the CCSD Z-vector method and with available experimental data.
@@ II.2.1 Permanent Dipole Moment (PDM)
The PDM operator () corresponding to a molecular system can be expressed as the sum of the electronic contribution and the nuclear contribution as,
In the above equation, e is the charge of the electron, is the position vector of electron from the origin, is the atomic number and is the position vector of nucleus. Now using Eq.(43) in Eq.(42) one can get the expectation value for the PDM, as
The electronic contribution to in Eq.(44) depends on the choice of correlation method and basis set, whereas the nuclear part is independent of both the correlation method and the basis set; it remains constant for a particular molecule with a certain charge and at a specific geometry.
@@ II.2.2 Hyperfine Structure (HFS) Constant
The magnetic hyperfine splitting in atoms arises from the interaction between the electromagnetic field generated by the electrons and the magnetic dipole moment of the nucleus. Within the magnetic dipole approximation, the vector potential at the position of the electron due to the nuclear magnetic moment arising from nucleus k,is given by
where is the position vector of the electron relative to the nucleus. In a relativistic framework, the corresponding magnetic hyperfine interaction for all electrons can be written as
where denotes the Dirac matrices of the electron . For a diatomic molecule, the magnetic hyperfine interaction can be separated into parallel and perpendicular components of the magnetic HFS constant. Along the molecular axes, the component of the expectation value of the corresponding HFS operator yields the parallel constant () while the components perpendicular to the molecular axes (x and y) give the perpendicular constants (), expressed as [59, 60]
Where denotes the nuclear spin quantum number and is the projection of total electronic angular momentum along the z-axis of the diatomic molecule.
@@ II.2.3 Electric Field Gradient (EFG)
The EFG is a traceless symmetric second-rank tensor that describes the spatial variation of the electric field around a nucleus due to the distribution of surrounding charges. The EFG provides valuable information about the local electronic structure and symmetry of atoms or molecules, particularly in cases where the environment of the nucleus is asymmetric. Within the principal axis coordinate system, the expectation value of the component of EFG at the position () of nucleus K is defined as
The defined above can be written as the sum of two contributions for a molecular system
The first term in the above equation is the electronic contribution, named , which depends on the ground state wave function. The second term, nuclear contribution , which arises from the other nuclei in the molecule, depends only on the position of that remaining nucleus and is independent of the electronic wave function.